Identify the slope and y-intercept for each question
12. y=2/3 x + 4
13. y= -3x+1
14. 3x+2y=12

Answer by math_helper(2461)

Identify the slope and y-intercept for each question
12. y=2/3 x + 4
Compare this to y=mx + b, slope is m, y-intercept is b:
slope = 2/3
y-intercept = 4


13. y= -3x+1

14. 3x+2y=12
This is standard form for a line Ax + By = C. Here you can first solve for y then you will have the slope-intercept form y = mx+b

There are also formulas relating m and b but they're not necessary to remember if you remember how to solve for y (which is far more important to know how to do). To illustrate, I will derive those formulas relating A,B,C to m and b:
Ax + By = C
By = C - Ax
y = -(A/B)x + C/B

Slope is -A/B, y-intercept is C/B (I will soon forget this fact and by the time I answer another problem like this I will be telling the student to solve for y :-) )